MALE FEMALE
Count 28 21
Mean 72.8 72.4
Standard Deviation 5.3 7.7

A researcher wanted to see whether there was a significant difference in resting pulse rates for men and women. The data she collected is summarized to the right. For both groups, the data appear to be unimodal and symmetric. Is this evidence that resting pulse rates for men are higher than for women? Test an appropriate hypothesis using a significance level of 0.10 (? = 0.10). You can assume all conditions and assumptions are met.

a. Write the Hypothesis for the test. H0: HA:

b. Find the P-value. P-value:_______________

c. Give the conclusion in context. (3 things needed)

d. create a 90% confidence interval for the difference in mean resting pulse rates for men and women and interpret your interval in the context of the problem.
Answer #1

A researcher wanted to see whether there was a significant difference in resting pulse rates for men and women. The data she collected is summarized to the right. For both groups, the data appear to be unimodal and symmetric. Is this evidence that resting pulse rates for men are higher than for women? Test an appropriate hypothesis using a significance level of 0.10 (? = 0.10)

a. Write the Hypothesis for the test.


\ H_0: \mu_1 = \mu_2 \ H_1: \mu_1 > \mu_2

Upper tail test

b. Find the P-value. P-value: 0.4152

t = \frac{\bar{x_1} - \bar{x_2}}{\sqrt{ \frac{(n_1 - 1) s_1^2 + (n_2 - 1) s_2^2} {(n_1 +n_2 -2)} *{(\frac{1}{n_1} + \frac{1}{n_2})}}}

Pooled-Variance t Test for the Difference Between Two Means

(assumes equal population variances)

Data

Hypothesized Difference

0

Level of Significance

0.1

Population 1 Sample

Sample Size

28

Sample Mean

72.8

Sample Standard Deviation

5.3

Population 2 Sample

Sample Size

21

Sample Mean

72.4

Sample Standard Deviation

7.7

Intermediate Calculations

Population 1 Sample Degrees of Freedom

27

Population 2 Sample Degrees of Freedom

20

Total Degrees of Freedom

47

Pooled Variance

41.3666

Standard Error

1.8567

Difference in Sample Means

0.4000

t Test Statistic

0.2154

Upper-Tail Test

Upper Critical Value

1.2998

p-Value

0.4152

Do not reject the null hypothesis

c. Give the conclusion in context. (3 things needed)

since P value =0.4152 which is > 0.10 level of significance. Ho is not rejected.

There is not enough evidence that resting pulse rates for men are higher than for women.

d. create a 90% confidence interval for the difference in mean resting pulse rates for men and women and interpret your interval in the context of the problem.

CI = ( \bar x_1 - \bar x_2) \pm t*se

Confidence Interval Estimate

for the Difference Between Two Means

Data

Confidence Level

90%

Intermediate Calculations

Degrees of Freedom

47

t Value

1.6779

Interval Half Width

3.1154

Confidence Interval

Interval Lower Limit

-2.7154

Interval Upper Limit

3.5154

90% confidence interval for the difference in mean resting pulse rates for men and women is ( -2.7154, 3.5154).

We are 90% confident that true difference in mean resting pulse rates for men and women falls in the interval           ( -2.7154, 3.5154).